OpenStudy (anonymous):

x^3 - i + sqrt3 = 0. Solve for x.

OpenStudy (perl):

the equation is x^3 -i + sqrt(3) = 0 does it say how many solutions it must have?

OpenStudy (perl):

x^3 = i - sqrt(3) x = cubrt ( i - sqrt(3))

OpenStudy (anonymous):

@perl No it doesn't specify how many.

OpenStudy (perl):

the cube root of a complex number has three solutions, thats why i was wondering

OpenStudy (anonymous):

@perl Is that the final answer?

OpenStudy (perl):

so we want x = (-sqrt(3) + i ) ^(1/3) lets use demoivre's theorem to get all three solutions

OpenStudy (perl):

first change -sqrt(3) + i to polar trig form

OpenStudy (perl):

so r = sqrt( (-sqrt(3))^2 + 1^2 ) = sqrt( 3 + 1 ) = 2 theta = atan ( -sqrt(3), 1 )

OpenStudy (perl):

people also use arctan, i prefer the atan function since it tells you the exact angle

OpenStudy (perl):

the angle is 5pi/6 http://www.wolframalpha.com/input/?i=atan%28-sqrt%283%29%2C1%29

OpenStudy (perl):

so we have so far x = [2 ( cos (5pi/6) + i sin(5pi/6) ] ^(1/3)

OpenStudy (perl):

x = [2 ( cos (5pi/6 + 2pi*n) + i sin(5pi/6 + 2pi*n) ) ] ^(1/3) now use demoivres theorem x = [ 2^(1/3) ( cos ( 5pi/6* 1/3 + 2pi/3 * n ) + i ( sin5pi/6 * 1/3 + 2pi/3*n ) ]

OpenStudy (perl):

let n = 0,1,2

OpenStudy (perl):

this gives us 3 solutions x = [2^(1/3)(cos( 5pi/18 + 2pi/3* 0) + i( sin5pi/18 + 2pi/3*0 ))] x= [2^(1/3)(cos( 5pi/18 + 2pi/3*1) + i( sin5pi/18 + 2pi/3*1 ))] x = [2^(1/3)(cos( 5pi/18 + 2pi/3* 2) + i( sin5pi/18 + 2pi/3*2 ))]

OpenStudy (perl):

thats a long expression, which we can estimate

OpenStudy (perl):

x = .80986 + .9651555 i x = -1.24078 + .21878 i x = .430918 - 1.1839

OpenStudy (perl):

which agrees with wolfram http://www.wolframalpha.com/input/?i=solve+x^3+-+i+%2B+sqrt3+%3D+0